If 40% of a given number is 8, then what is 15% of the given number?
Answer by jim_thompson5910(35256)

Let x = the unknown number (ie the "given number")

40% of a given number is 8

40% of a given number is 8

0.4 * x = 8

0.4 x = 8

0.4x = 8

x = 8/0.4

x = 20

So the "given number" is 20

-------------------------------------------------------------------
Now take 15% of that given number:


20*0.15 = 3


So 15% of 20 is 3

=======================================================================

Answer:

15% of the given number is 3